OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6 HELP PLEASE

OpenStudy (anonymous):

\[3(x + 1) = -2(x - 1) + 6\]\[\huge Distribute!\]\[3x + 3 = -2x + 2 + 6\]Can you solve from here?

OpenStudy (lgbasallote):

3(x+1) = -2(x-1) + 6 use distributive property 3x + 3 = -2x + 2 + 6 combine like terms 3x + 3 = -2x + 8 add 2x to both sides 3x + 2x + 3 = 8 subtract 3 from both sides 3x + 2x = 8 - 3 now simplify

OpenStudy (theviper):

=>3x+3=-2x+2+6 =>3x+2x=2+6-3 =>5x=4 =>x=??

OpenStudy (theviper):

@Calcmathlete that's why I gave u a medal ;)

OpenStudy (anonymous):

Simplifying my above statement further: \[3x + 3 = -2x + 2 + 6 \implies 3x + 2x = 2 + 6 - 3 \implies 5x = 5 \implies ?\]

OpenStudy (anonymous):

is it 4?

OpenStudy (lgbasallote):

no

OpenStudy (lgbasallote):

how?

OpenStudy (theviper):

5x=5 x=5/5=1

OpenStudy (lgbasallote):

would've been fun if @michaelr35 discovered that....

OpenStudy (anonymous):

No... \[5x = 5\]\[\frac{5x}{5} = \frac 55\]\[\frac{\cancel{5}x}{\cancel{5}} = \frac{\cancel{5}}{\cancel{5}}\]\[x = ?

OpenStudy (anonymous):

Thanks you guys

OpenStudy (anonymous):

yall smart !!!!!

OpenStudy (anonymous):

\(\LaTeX \space Fail...\)